The owner of Maumee Ford-Volvo wants to study the relationship between the age of a car and its selling price. Listed below is a random sample of 12 used cars sold at the dealership during the last year.

Car Age (years) Selling Price ($000)
1 11 12.1
2 8 10.5
3 14 5.7
4 17 4.9
5 9 5.0
6 8 13.4
7 10 10.5
8 14 9.0
9 13 9.0
10 17 4.5
11 6 12.5
12 6 11.5

1.

Determine the regression equation. (Negative amounts should be indicated by a minus sign. Round your answers to 3 decimal places.)

a=

b=

2.

Estimate the selling price of an 7-year-old car (in $000). (Round your answer to 3 decimal places.)

Selling Price:_______

3.

Interpret the regression equation (in dollars). (Round your answer to the nearest dollar amount.)

For each additional year, the car decreases    in value.
Answer #1

1.

The estimated regression line is given by,

ŷ = Bo+B1x

ŷ = 15.87+B1* (-0.61)

2.

The estimated selling price for 7 years ols=d is:

ŷ = 15.87 +7*(-0.61)

ŷ =11.56

EStimated selling price of an 7-year-old car is 11.56

3.

Interpretation of regression equation:

For each additional year, the car decreases -0.61 in value.